OpenStudy (anonymous):

Can you solve ((e^y+e^(-y))/2)*((e^y+e^(-y))/2)

OpenStudy (anonymous):

Solve what? This is just an expression, do you want to simplify it or find its zeros?

OpenStudy (anonymous):

I think you want to simplify it. Actually the given expression doesn't have any zeros.

OpenStudy (anonymous):

I have to integrate it from 0- ln(2). I didn't know if you could simplify the expression.

OpenStudy (anonymous):

0 to ln(2) I should say.

OpenStudy (anonymous):

The given expression is equal to \(\cosh^2(x)\).

OpenStudy (anonymous):

With h standing for..?

OpenStudy (anonymous):

Oh you don't know the hyperbolic cosine function. It's better that you read about it. http://en.wikipedia.org/wiki/Hyperbolic_function

OpenStudy (anonymous):

You still can solve it without knowing that though. \[\int\limits ({e^y+e^{-y} \over 2})^2 dy=\int \frac{1}{4}(e^{2y}+e^{-2y}+2)dy=\cdots\]